OpenStudy (anonymous):

solve for x: x^2+11x+10=0

OpenStudy (anonymous):

(x+1)(x+10) x=-10 x=-1

OpenStudy (aripotta):

how'd you get that. i'm feeling incredibly stupid today. maybe because it's 6 am and i haven't slept

OpenStudy (anonymous):

thank you for the answers, I am running out of time on this quiz and I have a few more to go.

OpenStudy (anonymous):

i factorised it and gave the x values when its 0, (x+1)=0 when x+1=0, x=-1. same for the other one.

OpenStudy (aripotta):

but how did you factorize it? like, how did you know it was going to be +1 and +10...

OpenStudy (wwe123):

x^2+11x+10=0 x^2+10x+x+10 x(x+10)+1(x+10) (x+1)(x+10) x=-1 or x=-10